OpenStudy (anonymous):

A water tank that holds 60 L of water can be emptied in 24 min. How long will it take to empty a water tank that holds 280 L of water?

OpenStudy (anonymous):

In other words, set up a ratio and solve for x.

OpenStudy (texaschic101):

60/24 = 280/x (60 l to 24 min. = 280l to x minutes)--cross multiply (60)(x) = (280)(24) now solve for x

OpenStudy (anonymous):

thats not the answer if you set it up 60/x = 280/24

OpenStudy (texaschic101):

what is the answer ?

OpenStudy (anonymous):

112

OpenStudy (texaschic101):

That is not how I wrote it....I wrote 60/24 = 280/x...cross multiply... (60)(x) = (280)(24) 60x = 6720 (now divide by 60 on both sides to get x by itself) 60x/60 = 6720/60 x = 112

OpenStudy (texaschic101):

do you understand ?

OpenStudy (anonymous):

yes i do

OpenStudy (texaschic101):

good....glad to help :)